Today is January 1, 2009. The state of Iowa has offered your firm a subsidized loan. It will be in the amount of $10,000,000 at an interest rate of 5 percent and have ANNUAL (amortizing) payments over 3 years. The first payment is due today and your taxes are due January 1 of each year on the previous year's income. The yield to maturity on your firm's existing debt is 8 percent. What is the APV of this subsidized loan? If you rounded in your intermediate steps, the answer may be slightly different from what you got. Choose the closest. Multiple Choice ?$3,497,224.43 $417,201.05 $840,797
Answer #1
Amount (1) Annual Instalment(2) Interest(3)=0.05*1 Principal (4)=2-3 RemainingBalance (5)=1-4 PV@0.08(6) (4)*PV@0.08(7)
10000000 3880330.6 500000 3380330.6 6619669.4 0.926 3130186.14
6619669.4 3880330.6 330983.47 3549347.13 3070322.27 0.857 3041790.49
3070322.27 3880330.6 153516.1135 3726814.49 -656492.22 0.7938 2958345.34 9130321.97

ANuual Instalment=10000000/pvA(0.08,3)

10000000/2.5771(See present value annuity table)=3880330.6

Total (3130186.14+3041790.49+2958345.34)=9130321.97

nearest is 840797.